A car moving with a constant speed of 55.0km/hr completes a circular track in 7.50 minutes. Calculate the magnitude of the acceleration of the car.


1
Expert's answer
2021-05-10T09:20:42-0400

To be given in quest ion

Car speed=55km/h="\frac{275}{18}m/sec"

Time (t)=7.50 minutes=450sec

Circuler motion

"a_{c}=\frac{v^2}{r}\rightarrow(1)"

"T=\frac{2\pi r}{v}\rightarrow(2)"

eqution (1);(2) we can written as

"a_{c}=\frac{v^2}{{\frac{vT}{2\pi}}}\rightarrow(3)"

"a_{c}=\frac{2πv}{T}\rightarrow(4)"

Put value

"a_{c}=\frac{2π\times\frac{275}{18}}{450}=0.213m/sec^2"